I'm having a problem setting up a word problem. The problem is Jim is 4 times as old as his daughter is now, 7 years ago he was 11 times her age. How old is Jim? Could you hep me set it up to solve?
Answer by ewatrrr(24785)


Hi
Jim is 4 times as old as his daughter is now
Let x and 4x represent the daughter's and Jim's age NOW
Question states*** 7yrs ago...
4x-7 = 11(x-7)
Solving for x
70 = 7x
x = 10, duaghter's age NOW. Jim is now 40.
CHECKING our Answer***7yrs ago...
33 = 11*3